NOAA Office of Satellite and Product Operations

SATOPS News Report


Friday, September 19, 2014

Customers or users who want to call to ask about problems or report outages can use the following contacts:

– For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-683-1400

– To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

 

DOY 262



GOES
All GOES operations were nominal over the past 24 hours.
No anomalies affecting product delivery via GVAR or s/c health &
safety. For more GOES information please see:
http://www.ospo.noaa.gov/Operations/GOES/index.html


POES
All POES operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For more POES information please see:
http://www.ospo.noaa.gov/Operations/POES/index.html


DMSP
OPS53 rev#25358 POGO-B 14:25z Failed: SCU problem at POGO-B, Failed over to BOSS-C, MIR sent.

OPS53 rev#25365 COOK-B 02:03z LOST: Could not get connected using DDC1 or DDC2, Cook-B Links where down Smooth Data Lost Cat-1, PB Sec = 117 SC Sec = 3440, MIR Submitted ADO Notified (Left message)

OPS53 rev#25365 POGO-B 0213z Nominal: Tried to get Smooth data. Lost CH8, than could not get CMDs out due to obscura.

All other DMSP operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://ngdc.noaa.gov/eog/dmsp.html


JASON-2
All Jason-2 operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.


SNPP
SNPP rev#14985 14:06z Nominal: Weekly S/C Load, PROC hung at the end (all loads completed) and suspended but unable to abort in the central commanding window. Engineers and factory investigating the problem, CONTINUED support using manual window versus nominal central window, IN the back orbit recycled RP0, issue resolved.

All other SNPP operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For more SNPP Information please see:
http://npp.gsfc.nasa.gov/


ESPC reported data outages or delays

Data Lost:
ESPC was informed by FNMOC/Monterey that because of problems with the data, DMSP F-18 rev 25365 is unrecoverable and considered lost.
ESPC did not receive the 0000 and 0030 UTC Meteosat-7 data due to a problem at EUMetSat.



Data Delay:
EWA beach data still not being received, Initial outage 15 Sept, EWA beach still investigating.
1937Z - EWA BEACH DATA NOW BEING RECEIVED
1653Z - SNPP data outage/delay all AFWA subscriptions suspended

The Data Distribution System (DDS) began having large backlogs of file transfers to all AFWA master destinations at approximately 15:55 UTC. Engineering was notified and are troubleshooting with AFWA personnel. All AFWA subscriptions were suspended at 16:53 UTC and will remain suspended until the reason for the DDS failure is determined. ESPC did not receive the 1500Z, 1600Z and 1700Z North and South CLOUD data at OFFUTT/AFWA. The issue is being investigated.

September 19, 2014 0200 UTC - ESPC reports that all the backlogged cloud data from AFWA has been received.

2200Z - Ewa Beach data for N18 and N19 still not processing. MetOp data is being received.
0100Z - Users notified IDPS that subscriptions for AFWA data had been restored and data began processing at 0010Z.